# Making Parametric Anomaly Detection on Tabular Data Non-Parametric Again

### v1

@@ Abstract.
Deep learning for tabular data has garnered increasing attention in recent years, yet employing deep models for structured data remains challenging. While these models excel with unstructured data, their efficacy with structured data has been limited. Recent research has introduced retrieval-augmented models to address this gap, demonstrating promising results in supervised tasks such as classification and regression. In this work, we investigate using retrieval-augmented models for anomaly detection on tabular data. We propose a reconstruction-based approach in which a transformer model learns to reconstruct masked features of normal samples. We test the effectiveness of KNN-based and attention-based modules to select relevant samples to help in the reconstruction process of the target sample. Our experiments on a benchmark of 31 tabular datasets reveal that augmenting this reconstruction-based anomaly detection (AD) method with non-parametric relationships via retrieval modules may significantly boost performance. Our code to reproduce the experiments is made available on GitHub.
Anomaly Detection, Tabular Data, Deep Learning
@@ 1. Introduction
Supervised anomaly detection (AD) consists in learning to characterize a normal distribution using a dataset only composed of samples belonging to the normal [^1] The term normal relates to the concept of normality, in opposition to anomaly. class, in order to identify in a separate dataset the samples that do not belong to this normal distribution, namely anomalies. This type of approach is often used when the imbalance between classes is too severe, causing standard supervised approaches to fail (Yanmin et al., 2011). Examples of such applications are cyber intrusion detection (Ben Said et al., 2023), fraud detection on credit card payment (Hilal et al., 2022; Thimonier et al., 2023b), or tumor detection on images (Wei et al., 2018). On the contrary, Unsupervised anomaly detection refers to identifying anomalies in a dataset without using labeled training data. These algorithms aim to discover patterns or structures in the data and flag instances that deviate significantly from these patterns. The application of such an approach usually includes detecting mislabeled samples or removing outliers from a dataset that may hinder the training of a model.
@@
While deep learning methods have become ubiquitous and are widely used in the industry for various tasks on unstructured data, relying on deep models for tabular data remains challenging. Indeed, Grinsztajn et al. (2022) discuss how the inherent characteristics of tabular data make this type of data challenging to handle by standard deep models. Hence, recent research on deep learning for structured data (Shavitt and Segal, 2018; Arik and Pfister, 2021; Somepalli et al., 2021; Kossen et al., 2021; Gorishniy et al., 2023) has been oriented towards proposing novel training frameworks, regularization techniques or architectures, tailored for tabular data. Similarly, the best-performing anomaly detection algorithms proposed in the literature involve accounting for the particular structure of tabular data while general approaches tend to struggle. For instance, in (Bergman and Hoshen, 2020; Qiu et al., 2021; Shenkar and Wolf, 2022; Thimonier et al., 2023a) authors put forward self-supervised anomaly detection algorithms targeted for tabular data.
In particular, recent research has emphasized the pivotal role of non-parametric relations between samples in fostering deep learning model’s performance (Gorishniy et al., 2023; Kossen et al., 2021; Somepalli et al., 2021; Thimonier et al., 2023a). Kossen et al. (2021) and Somepalli et al. (2021) propose novel deep architecture with internal retrieval models that allow leveraging inter-sample relations. Similarly, Gorishniy et al. (2023) investigate the pertinence of including non-parametric mechanisms by augmenting an MLP with external retrieval modules and show a significant increase in performance for classification and regression tasks on tabular data. Parallel to this line of work, the reconstruction-based AD method of Thimonier et al. (2023a) relies on non-parametric transformers (Kossen et al., 2021) to leverage sample-sample dependencies and shows significant improvement in anomaly detection accuracy in comparison with existing methods.
@@
As empirically demonstrated in (Thimonier et al., 2023a), both feature-feature and sample-sample dependencies should be conjointly learned by a model to detect anomalies on tabular data efficiently. Indeed, Han et al. (2022) discusses how anomalies on tabular data can be categorized into $4$ types of anomalies. We believe leveraging both interactions is necessary to efficiently identify all $4$ types. First, dependency anomalies, referring to samples that do not follow the dependency structure that normal data follow, require feature-feature dependencies to be efficiently identified. Second, global anomalies refer to unusual data points that deviate significantly from the norm. Relying on both dependencies should improve a model’s capacity to detect these anomalies. Third, local anomalies that refer to the anomalies that are deviant from their local neighborhood can only be identified by relying on sample-sample dependencies. Finally, clustered anomalies, also known as group anomalies, are composed of anomalies that exhibit similar characteristics. This type of anomaly requires feature-feature dependencies to be identified.
@@
Hence, we further explore the possibility of leveraging both feature-feature and sample-sample relations for supervised anomaly detection by investigating how existing AD methods can be augmented by an external retrieval module. External retrieval modules can be considered as instrumental as they can augment any existing model. In contrast, most models that rely on internal retrieval mechanisms (e.g. non-parametric transformers (Kossen et al., 2021)) are bound to some inductive biases and cannot be used to learn all possible tasks that may be relevant for anomaly detection. We test the relevance of external modules by employing transformers (Vaswani et al., 2017) in a mask-reconstruction framework to construct an anomaly score since it was proven to offer strong anomaly detection performance (Thimonier et al., 2023a). We implement several external retrieval methods to augment the vanilla transformer and evaluate the performance of each approach on an extensive benchmark of tabular datasets. We empirically show that the tested approaches incorporating retrieval modules to account for the sample-sample relations outperform the vanilla transformer that only attends to feature-feature dependencies.
@@ 2. Related works
Ruff et al. (2021) discuss how anomaly detection bears several denominations that more or less designate the same class of algorithms: anomaly detection, novelty detection, and outlier detection. Overall, the literature is comprised of $4$ main classes of anomaly detection algorithms: density estimation, one-class classification, reconstruction-based and self-supervised algorithms.
@@
Other anomaly detection methods focus on learning to reconstruct samples belonging to the normal distribution. In inference, the capacity of the model to reconstruct an unseen sample is used as a measure of anomalousness. The more capable a model is to reconstruct a sample, the more likely the sample is to belong to the normal distribution seen in training. Such approach include methods involving autoencoders (Chen and Konukoglu, 2018; Kim et al., 2020), diffusion models (Wolleb et al., 2022; Zhang et al., 2023), GANs (Schlegl et al., 2017) or attention-based models (Thimonier et al., 2023a).
@@
One-class classification describes the task of identifying anomalies without directly estimating the normal density. This class of algorithm involves discriminative models that directly estimate a decision boundary. In (Schölkopf et al., 1999; Tax and Duin, 2004; Ruff et al., 2018), the authors propose algorithms that estimate the support of the normal distribution, either in the original data space or in a latent space. During inference, one flags the samples outside the estimated support as anomalies. Other one-class classification methods include tree-based approaches such as isolation forest (IForest) (Liu et al., 2008), extended isolation forest (Hariri et al., 2021), RRCF (Guha et al., 2016) and PIDForest (Gopalan et al., 2019). Other methods also include approaches to augment existing one-class classification methods with a classifier through generating synthetic anomalies during training, such as DROCC (Goyal et al., 2020).
@@
Given the recent successes of self-supervision for a large variety of tasks, researchers have also investigated using self-supervised methods for anomaly detection. Bergman and Hoshen (2020) and Qiu et al. (2021) propose transformation based anomaly detection methods for tabular data. The former relies on a classifier’s capacity to identify which transformation was applied to a sample to measure anomalousness, while the latter relies on a contrastive approach. Similarly, Shenkar and Wolf (2022) also propose a contrastive approach to flag anomalies by learning feature-feature relation for normal samples. Parallel to this line of work, (Sohn et al., 2021; Reiss and Hoshen, 2023) have focused on proposing self-supervised approaches for representation learning tailored for anomaly detection.
@@
Retrieval modules have gained attention in recent years in many fields of machine learning. For instance, Blattmann et al. (2022) introduces a retrieval module to foster the scalability and efficiency of diffusion models. Parallel to that, Lin et al. (2022) introduced retrieval for cross-task generalization of large language models, and Chen et al. (2022) introduced it to enhance prompt learning. Finally, retrieval methods have been increasingly used to increase the performance of deep models for tabular data. For instance, Kossen et al. (2021) and Somepalli et al. (2021) introduced internal retrieval modules in deep architecture for supervised tasks on tabular data, while Thimonier et al. (2023a) relied on internal retrieval modules for anomaly detection. Finally, Gorishniy et al. (2023) investigated using external retrieval modules to augment an MLP for supervised tasks on tabular data.
@@
Figure 1. Forward pass for sample $\mathbf{z}$, see appendix A.2.1 for more detail on training procedure. In the case of no retrieval module, the prediction for a sample $\mathbf{z}$ consists of the upper part of the figure with $\lambda=0$.
@@ 3.1. Learning Objective
Let $\mathcal{D}_{train}=\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$ represent the training set composed of $n$ normal samples with $d$ features. The standard approach to anomaly detection involves learning some function $\phi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{p}$ by minimizing a loss function $\mathcal{L}$. The chosen loss function will vary according to the class of the considered anomaly detection algorithm. Nevertheless, the overall aim of $\mathcal{L}$ is to characterize the distribution of the samples in the training set as precisely as possible. Depending on the chosen AD algorithm, the obtained representation $\phi_{\theta}(\mathbf{x})$ of sample $\mathbf{x}$ can be used directly or indirectly to compute an anomaly score.
@@
where $\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x}))$ will vary according to the chosen task. For a reconstruction-based method $\mathcal{L}$ can be the squared $\ell_{2}$-norm of the difference between the original sample $\mathbf{x}$ and its reconstructed counterpart, $\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x}))=\|\mathbf{x}-\phi_{\theta}(\mathbf{x})\|^{2}$.
Introducing an external retrieval module permits keeping the original objective unchanged while augmenting $\phi_{\theta}$ with non-parametric mechanisms. The model involves non-parametric relations because it leverages the entire training dataset when making its prediction. Hence, to optimize its objective the model can conjointly attend to feature-feature and sample-sample interactions.
Formally, instead of minimizing the loss as described in equation 1, the parameters $\theta$ of function $\phi_{\theta}$ are optimized to minimize the loss function as follows
@@
Nevertheless, not all approaches to AD may benefit from such non-parametric mechanisms. Some pretext tasks involving sample-sample dependencies may lead to degenerate solutions, e.g. approaches based on contrastive learning such as the approaches of Qiu et al. (2021) or Shenkar and Wolf (2022). However, reconstruction-based AD methods appear as a natural class of algorithms that may benefit from these non-parametric mechanisms.
@@
We empirically investigate the pertinence of external retrieval modules, as detailed in section 3.2, on a similar method to Thimonier et al. (2023a) involving stochastic masking and mask reconstruction for tabular data. Stochastic masking consists in masking each entry in a sample vector $\mathbf{x}\in\mathbb{R}^{d}$ with probability $p_{mask}$, and the objective task is to predict the masked-out features from the unmasked features. Formally, let $m\in\mathbb{R}^{d}$ be a binary vector taking value $1$ when the corresponding entry in $\mathbf{x}$ is masked, $\mathbf{x}^{m}=\{x_{j}:m_{j}=1\}$ represents the masked entries of sample $\mathbf{x}$, and $\mathbf{x}^{o}=\{x_{j}:m_{j}=0\}$ its unmasked counterpart. We evaluate the benefit of replacing the traditional reconstruction learning objective defined as
(3) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}d(\mathbf{x}^{m},\phi_{\theta}(\mathbf{x}^{o})),$
@@
(4) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}d\left(\mathbf{x}^{m},\phi_{\theta}\left(\mathbf{x}^{o}\mid\mathcal{D}_{train}^{o}\right)\right),$
@@
Table 1. Comparison of transformer-based methods.
@@
F1-Score $56.2$ $56.1$ $55.7$ $\mathbf{58.6}$ $53.9$
AUROC $83.4$ $83.1$ $83.1$ $\mathbf{84.4}$ $82.1$

### v2

# Retrieval Augmented Deep Anomaly Detection for Tabular Data

@@ Abstract.
Deep learning for tabular data has garnered increasing attention in recent years, yet employing deep models for structured data remains challenging. While these models excel with unstructured data, their efficacy with structured data has been limited. Recent research has introduced retrieval-augmented models to address this gap, demonstrating promising results in supervised tasks such as classification and regression. In this work, we investigate using retrieval-augmented models for anomaly detection on tabular data. We propose a reconstruction-based approach in which a transformer model learns to reconstruct masked features of normal samples. We test the effectiveness of KNN-based and attention-based modules to select relevant samples to help in the reconstruction process of the target sample. Our experiments on a benchmark of 31 tabular datasets reveal that augmenting this reconstruction-based anomaly detection (AD) method with sample-sample dependencies via retrieval modules significantly boosts performance. The present work supports the idea that retrieval module are useful to augment any deep AD method to enhance anomaly detection on tabular data. Our code to reproduce the experiments is made available on GitHub.
Anomaly Detection; Tabular Data; Deep Learning
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679559ccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Computing methodologies Semi-supervised learning settingsccs: Computing methodologies Neural networks
@@ 1. Introduction
Semi-supervised anomaly detection (AD) consists in learning to characterize a normal distribution using a dataset only composed of samples belonging to the normal [^1] The term normal relates to the concept of normality, in opposition to anomaly. class, in order to identify in a separate dataset the samples that do not belong to this normal distribution, namely anomalies. This class of algorithms is often used when the imbalance between classes is too severe, causing standard supervised approaches to fail (Yanmin et al., 2011). Examples of such applications are cyber intrusion detection (Ben Said et al., 2023), fraud detection on credit card payment (Hilal et al., 2022; Thimonier et al., 2023), or tumor detection on images (Wei et al., 2018). On the contrary, unsupervised anomaly detection refers to identifying anomalies in a dataset without using labeled training data. These algorithms aim to discover patterns or structures in the data and flag instances that deviate significantly from these patterns. The application of such an approach usually includes detecting mislabeled samples or removing outliers from a dataset that may hinder a model’s training process.
@@
While deep learning methods have become ubiquitous and are widely used in the industry for various tasks on unstructured data, relying on deep models for tabular data remains challenging. Indeed, Grinsztajn et al. (2022) discuss how the inherent characteristics of tabular data make this type of data challenging to handle by standard deep models. Hence, recent research on deep learning for structured data (Shavitt and Segal, 2018; Arik and Pfister, 2021; Somepalli et al., 2021; Kossen et al., 2021; Gorishniy et al., 2023) has been oriented towards proposing novel training frameworks, regularization techniques or architectures tailored for tabular data. Similarly, general AD methods appear to struggle with tabular data, while the best-performing AD algorithms on tabular data involve accounting for the particular structure of this data type. For instance, (Bergman and Hoshen, 2020; Qiu et al., 2021; Shenkar and Wolf, 2022; Thimonier et al., 2024) put forward self-supervised anomaly detection algorithms targeted for tabular data that significantly outperform general methods on most tested datasets.
In particular, recent research has emphasized the pivotal role of combining feature-feature and samples-sample dependencies in fostering deep learning model’s performance on tabular data (Gorishniy et al., 2023; Kossen et al., 2021; Somepalli et al., 2021). Following these recent findings, we investigate the benefits of including external retrieval modules to leverage sample-sample dependencies to augment existing AD methods. External retrieval modules can be considered instrumental as they can augment any existing model that may only rely on feature-feature dependencies. In contrast, models that rely on internal retrieval mechanisms are bound to some inductive biases and cannot be used to learn all possible tasks that may be relevant for anomaly detection.
@@
Leveraging both types of dependencies is critical to detect all types of anomalies effectively and can increase consistency across datasets, as we empirically show in section 5.1. Han et al. (2022) categorize anomalies in tabular data into $4$ families of anomalies which require different types of dependencies to be correctly identified. First, dependency anomalies explicitly refers to samples that do not follow the dependency structure that normal data follow require feature-feature dependencies to be efficiently identified. Second, global anomalies refer to unusual data points that deviate significantly from the norm. Relying on both dependencies should improve a model’s capacity to detect these anomalies. Third, local anomalies that refer to the anomalies that are deviant from their local neighborhood can only be identified by relying on sample-sample dependencies. Finally, clustered anomalies, also known as group anomalies, are composed of anomalies that exhibit similar characteristics. This type of anomaly requires feature-feature dependencies to be identified.
@@
We test the relevance of external modules by employing transformers (Vaswani et al., 2017) in a mask-reconstruction framework to construct an anomaly score as it was proven to offer strong anomaly detection performance (Thimonier et al., 2024). We implement several external retrieval methods to augment the vanilla transformer and evaluate the performance of each approach on an extensive benchmark of tabular datasets.
We empirically show that the tested approaches incorporating retrieval modules to account for the sample-sample relations outperform the vanilla transformer that only attends to feature-feature dependencies. Furthermore, we propose an empirical experiment to account for the pertinence of combining dependencies, showing that detecting some types of anomalies can require a particular type of dependency.
@@
  • We compare our approach to existing methods found in the literature and observe that our method obtains competitive performance metrics.
  • We provide an explanation as to why combining dependencies leads to better identification of anomalies in tabular data.
@@ 2. Related works
Ruff et al. (2021) discuss how anomaly detection bears several denominations that more or less designate the same class of algorithms: anomaly detection, novelty detection, and outlier detection. The literature comprises $4$ main classes of anomaly detection algorithms: density estimation, one-class classification, reconstruction-based, and self-supervised algorithms.
@@
Other anomaly detection methods focus on learning to reconstruct samples belonging to the normal distribution. In inference, the capacity of the model to reconstruct an unseen sample is used as a measure of anomalousness. The more capable a model is to reconstruct a sample, the more likely the sample is to belong to the normal distribution seen in training. Such approach include methods involving autoencoders (Chen and Konukoglu, 2018; Kim et al., 2020), diffusion models (Wolleb et al., 2022; Zhang et al., 2023), GANs (Schlegl et al., 2017) or attention-based models (Thimonier et al., 2024).
@@
One-class classification describes the task of identifying anomalies without directly estimating the normal density. This class of algorithm involves discriminative models that directly estimate a decision boundary. In (Schölkopf et al., 1999; Tax and Duin, 2004; Ruff et al., 2018), the authors propose algorithms that estimate the support of the normal distribution, either in the original data space or in a latent space. During inference, one flags the samples outside the estimated support as anomalies. Other one-class classification methods include tree-based approaches such as isolation forest (IForest) (Liu et al., 2008), extended isolation forest (Hariri et al., 2021), RRCF (Guha et al., 2016) and PIDForest (Gopalan et al., 2019). Other methods include approaches to augment existing one-class classification methods with a classifier by generating synthetic anomalies during training, such as DROCC (Goyal et al., 2020).
@@
Given the recent successes of self-supervision for many tasks, researchers have also investigated using self-supervised methods for anomaly detection. (Bergman and Hoshen, 2020) and (Qiu et al., 2021) propose transformation based anomaly detection methods for tabular data. The former relies on a classifier’s capacity to identify which transformation was applied to a sample to measure anomalousness, while the latter relies on a contrastive approach. Similarly, (Shenkar and Wolf, 2022) also proposes a contrastive approach to flag anomalies by learning feature-feature relation for normal samples. Parallel to this line of work, (Sohn et al., 2021; Reiss and Hoshen, 2023) have focused on proposing self-supervised approaches for representation learning tailored for anomaly detection.
@@
Retrieval modules have gained attention in recent years in many fields of machine learning. For instance, (Blattmann et al., 2022) introduces a retrieval module to foster the scalability and efficiency of diffusion models. Parallel to that, (Lin et al., 2022) introduced retrieval for cross-task generalization of large language models, and (Chen et al., 2022) introduced it to enhance prompt learning. Finally, retrieval methods have been increasingly used to increase the performance of deep models for tabular data. For instance, (Kossen et al., 2021) and (Somepalli et al., 2021) introduced internal retrieval modules in deep architecture for supervised tasks on tabular data, while (Thimonier et al., 2024) relied on internal retrieval modules for anomaly detection. Finally, (Gorishniy et al., 2023) investigated using external retrieval modules to augment an MLP for supervised tasks on tabular data.
@@
Figure 1. Forward pass for sample $\mathbf{z}$, see section 3.4 for more detail on training procedure. In the case of no retrieval module, the prediction for a sample $\mathbf{z}$ consists of the upper part of the figure with $\lambda=0$.
@@ 3.1. Learning Objective
Let $\mathcal{D}_{train}=\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$ represent the training set composed of $n$ normal samples with $d$ features. The standard approach to anomaly detection involves learning some function $\phi_{\theta}:\mathbb{R}^{d}\to\mathcal{Z}$ by minimizing a loss function $\mathcal{L}$. The chosen loss function and the space $\mathcal{Z}$ will vary according to the class of the considered anomaly detection algorithm. Nevertheless, the overall aim of $\mathcal{L}$ is to characterize the distribution of the samples in the training set as precisely as possible. Depending on the chosen AD algorithm, the obtained representation $\phi_{\theta}(\mathbf{x})$ of sample $\mathbf{x}$ can be used directly or indirectly to compute an anomaly score.
@@
where $\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x}))$ will vary according to the chosen task. For a reconstruction-based method, $\mathcal{Z}$ can be the original data space and $\mathcal{L}$ can be the squared $\ell_{2}$-norm of the difference between the original sample $\mathbf{x}$ and its reconstructed counterpart, $\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x}))=\|\mathbf{x}-\phi_{\theta}(\mathbf{x})\|^{2}$.
Introducing an external retrieval module permits keeping the original objective unchanged while augmenting $\phi_{\theta}$ with sample-sample dependencies through non-parametric mechanisms. The model involves non-parametric relations as it leverages the entire training dataset to make its prediction. Hence, the model can conjointly attend to feature-feature and sample-sample interactions to optimize its objective.
Formally, instead of minimizing the loss as described in eq. (1), the parameters $\theta$ of the function $\phi_{\theta}$ are optimized to minimize the loss function as follows
@@
Nevertheless, not all approaches to AD may benefit from such non-parametric mechanisms. Some pretext tasks involving sample-sample dependencies may lead to degenerate solutions, e.g. approaches based on contrastive learning such as the approaches of (Qiu et al., 2021) or (Shenkar and Wolf, 2022). However, reconstruction-based AD methods appear as a natural class of algorithms that may benefit from these non-parametric mechanisms.
@@
In the mask reconstruction context, we empirically investigate the pertinence of external retrieval modules, as detailed in section 3.2. Our approach includes stochastic masking which consists in masking each entry in a sample vector $\mathbf{x}\in\mathbb{R}^{d}$ with probability $p_{mask}$ while setting as the objective task the prediction of the masked-out features from the unmasked features. Formally, we sample $\mathbf{m}\in\mathbb{R}^{d}$ a binary mask vector taking value $1$ when the corresponding entry in $\mathbf{x}$ is masked, 00 otherwise. This mask $\mathbf{m}$ is then used to construct $\mathbf{x}^{m},\mathbf{x}^{o}\in\mathbb{R}^{d}$ representing respectively the masked and unmasked entries of sample $\mathbf{x}$. $\mathbf{x}^{m},\mathbf{x}^{o}$ are obtained as follows,
(3) $\begin{array}[]{rl}\mathbf{x}^{m}&=\mathbf{m}\odot\mathbf{x}\\ \mathbf{x}^{o}&=(\mathbf{1}_{d}-\mathbf{m})\odot\mathbf{x},\end{array}$
where $\mathbf{1}_{d}$ is the $d$-dimensional unit vector.
(a) F1-score ($\uparrow$)
(b) Rank ($\downarrow$)
Figure 2. For each of the $31$ datasets on which models were evaluated, we report the average F1-score over $20$ runs for $20$ different seeds. We refer readers to Thimonier et al. (2024) for details on the obtained metrics and the hyperparameters used for each method. For both figures, the model displayed on the far left is the worst-performing model for the chosen metric, and the one on the far right is the best-performing model. We also highlight the metric of the best-performing model in bold.
A model $\phi_{\theta}:\mathbb{R}^{d}\times\{0,1\}^{d}\to\mathbb{R}^{d}$ is trained to reconstruct the mask features $\mathbf{x}^{m}$ from its unmasked counterpart $\mathbf{x}^{o}$ and the mask vector $\mathbf{m}$. By construction, $\phi_{\theta}(\mathbf{x}^{o};\mathbf{m})$ only has non-zero values for corresponding masked features in $\mathbf{m}$.
In the present work, we evaluate the benefit of replacing the traditional reconstruction learning objective defined as
(4) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}d(\mathbf{x}^{m},\phi_{\theta}(\mathbf{x}^{o};\mathbf{m})),$
@@
(5) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}d\left(\mathbf{x}^{m},\phi_{\theta}\left(\mathbf{x}^{o};\mathbf{m},\mathcal{D}_{train}^{o}\right)\right),$
@@
Table 1. Comparison of transformer-based methods. We observe that the external retrieval module attention-bsim significantly improves the AD performance of the vanilla transformer by $4.3\%$ regarding the F1-Score and $1.2\%$ for the AUROC.
@@
F1-Score ($\uparrow$) $56.2$ $56.1$ $55.7$ $\mathbf{58.6}$ $53.9$
AUROC ($\uparrow$) $83.4$ $83.1$ $83.1$ $\mathbf{84.4}$ $82.1$